find a so that the graph of f(x) =LOGA ^ XCONTAINS THE POINT (21,4)
A=

Answer by stanbon(48535)

find a so that the graph of f(x) =LOGA ^ XCONTAINS THE POINT (21,4)
y=log(base a)x
This means:
x=a^y
Substituting (21,4) you get:
21=a^4
a= 21^(1/4)=2.140695143...
Cheers,
Stan H.